Find the value of c that completes the square.
y to the second power - 3y+ c

Answer by checkley71(8403)

Y^2-3Y+C
(Y-3/2)(Y-3/2) THEREFORE C=-3/2*-3/2=9/4
Y^2-3Y+9/4